OpenStudy (anonymous):

Question: A football team lost 9 yards on its first play. It then gained 9 yards on its second play. A. What expression can be written to represent how many yards the football team is from where it started? Explain your reasoning. B. How many yards is the football team from where it started? C. Suppose that, instead, the football team had gained 9 yards on its first play and lost 9 yards on its second play. Would the team’s position on the field have been any different than it actually was after the first two plays? Why or why not?

OpenStudy (ccswims):

Ok, I'm pretty sure that A. is -9 + 9, I don't see why it would need any variables

OpenStudy (ccswims):

B. It is currently at 0 yards, assuming that the were originally at 0 yards it did not change

OpenStudy (ccswims):

C. It will not change because -9 + 9 = 0 and 9 - 9 = 0

OpenStudy (anonymous):

Thank you!!!!!:D

OpenStudy (anonymous):

but how do you know how much they had at the start thats why i think u would use a variable

OpenStudy (ccswims):

idk (x-9) (x + 9) This seems so simple but idk not today

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@aum @arabpride

OpenStudy (arabpride):

All they did was gain what they lost therefore, it's like they're standing where they began~ SO at zero... B. Zero yards because they ended up where they started~ C. They remain at point "zero" because the only difference here was that the process was reversed, but the results remained the same~

OpenStudy (aum):

A. (-9 + 9) is the expression. Loss in yards is expressed as negative yards and gain in yards is expressed as plus yards. B. -9 + 9 = 0. Zero yards from the starting position. C. (+9 - 9) is the expression. It is equal to zero which is same as the answer for A. The reason they are the same is because -9 + 9 = 9 - 9 = 0 and this is the commutative property of addition.

OpenStudy (anonymous):

Thank you!!!!